# Evaluating from Benign to Dynamic Adversarial: A Squid Game for Large Language Models

### v1

@@ 3.1 Evaluation Setup
Regarding the red-green light game, we empirically set the probability of an interruption instruction $\rho$ to $0.4$. We sample 50 items from the instruction-following and reasoning sets of the LiveBench (White et al., 2025) to construct a long task set. For sugar honeycombs, we collect metadata (e.g., task description, canonical solution, and unit tests) from two famous datasets for code evaluation, HumanEval (Chen et al., 2021) and LiveCodeBench (Jain et al., 2025), and utilize the abstract syntax tree to build degraded codes where redundant logic or irrelevant contents were injected. Specifically, we include $5$ strict unit tests, such as input, consistency (computed by CodeBERTScore (Zhou et al., 2023)), for each $C_{out}$. For tug of war, we collect debating topics, as well as the pro and con opinions, from the latest WUDC 2025111https://wudc2025.calicotab.com/open/. In marbles, we adopt a free-form question setting in which a model generates its own questions and reference answers to challenge another model. For glass stepping stones, we first generate a safety route and then instruct models on order selection after releasing the game details. In the final game, we collect safety-related QA pairs from SaladBench (Li et al., 2024b) and guide the attacker to modify them to be harmful inputs to the defender. Each attacker-defender pair undergoes a role swap to ensure fairness. See more details in App. C.
@@ 3.2 Main Results
Model Performance w.r.t. Game Settings. In Tab. 1, we provide the overall statistics of six levels in the Squid Game. It can be observed that the elimination rate in red-green light game significantly surpasses all games except the final round. This mirrors the bottleneck of the existing LLMs in strict instruction following and interruption handling perspectives. In another static evaluation scenario (sugar honeycombs), the survival situation is much better, but the standard deviation is much higher than the first game, showing relatively large performance gaps between the remaining models across different rounds. This phenomenon is also evident in the dynamic evaluation environment of marbles. In comparison, the glass stepping stones game owns the lowest survival rate in dynamic adversarial scenarios except the final round, indicating that incremental reasoning requirement still challenging LLMs. As exhibited in Fig. 3, Gemini 2.5 Pro, GPT-5 and a range of reasoning LLMs exhibit comparatively superior performance, with the improvement most pronounced in the Qwen family, where the reasoning-enabled Qwen3-8B surpasses its 72B predecessor, Qwen2.5-72B.
@@
Most LLMs Fail to React Promptly. Fig. 4 shows the elimination points of participating LLMs in the red-green light game, where interruption are randomly introduced to a long task for instruction-following and robustness evaluation. We observe that Gemini 2.5 and Seed 1.6 series, as well as OpenAI’s reasoning models achieve a near-perfect pass rate, while most LLMs struggle to return correct security code or occur context discontinuity. It is worth noting that none of the early version of GPTs and Claudes (e.g., GPT-4.1 and Claude 3.7 Sonnet) reach the finish line, which is also observed for the recent GLM 4.5. This can largely be attributed to their poor math abilities in constructing the security codes. We also find lightweight LLMs, such as Mini, Flash, Air versions, perform better than their standard version, showing the advantage of lightweighting in this regard.
@@
Some Models Exploit Loopholes in the Game. In the red-green light game, we notice that some weaker models, such as Qwen 2.5, Llama 3.x, and DeepSeek V3 series, managed to pass the game by exploiting its loopholes. For example, these models simply modify the numerical sequence of the given security code to achieve the target sum, indicating the potential of evaluation methodology leakage in the current static benchmarks.
@@
Open-Source Models Lag Behind on Sugar Honeycombs. In Fig. 5, we show the testing results of code modified by different LLMs with CodeBERTScore reported. First, the sampled degraded code have similar CodeBERTScore to ensure fairness. We observe that the code generated by LLMs with limited reasoning capabilities often failed to satisfy all test requirements. Specifically, although exhibiting high test pass rates, the CodeBERTScore of Gemini 2.5 series is significantly lower than other models. We attribute this observation to two factors, i.e., a large volume of superfluous comments and verbose naming for variables and functions.
Figure 6: Statistics of the quotas used by different teams. “RiiMjj” is the $j$-th match in $i$-th round.
Figure 7: Left. Model combat style analysis: offensive vs. defensive success rates. The red dotted circle encompasses other models with extreme defensive behavior, covering nearly all top proprietary LLMs. Right. Word cloud of all spontaneously generated questions during the marbles game.
@@
Resource Consumption Comparison in Adversarial Dialogues. We visualize the team formation of the tug of war game as a graph in Fig. 9 and report the detailed matchup in Fig. 10. Fig. 6 exhibits the quota usage of teams in different matches. It is obvious that the team composed of lightweight models consumes noticeably fewer quota than those with heavily-reasoning LLMs. Surprisingly, across the five matches in which teams were eliminated owing to exhausted resources, DeepSeek-R1-0528 figured in 60 % of the cases, while GPT-5 and Claude-4-Opus-Thinking account for the rest of 40%. This highlights the necessity of improving the utilization rate of output tokens, especially in resource-constrained scenarios. Additionally, we observe a clear shift in standpoint via adversarial dialogues, demonstrating the effectiveness of our design to evaluate such a security flaw. Since the reversal of the logical structure of the argument is particularly evident after rounds of persuasive dialogues, the agreement rate between the LLM judge and humans is over 95%.
@@
Autonomously Generated Questions Struggle to Differentiate Top Models. In the marbles game, we instruct the models to freely set questions to stump each other as much as possible. Fig. 7 exhibits the combat success results and the word cloud for autonomously generated questions. It can be seen that almost all LLMs perform a higher defensive success rate than offensive (landing within the gray quarter-circle), indicating that the completely autonomous questioning strategy struggles to challenge top models. The word cloud reveals that the models primarily generate knowledge-based questions.
@@ 3.3 Comparison to Other LLM Benchmarks
We further compare Squid Game to three commonly used benchmarks, ChatBot Arena (Chiang et al., 2024), LiveBench (White et al., 2025), and LiveCodeBench (Jain et al., 2025). In Fig. 8, we provide scatter plots and correlation coefficients for models that have been evaluated on both benchmarks. First, we compare the total number of steps each model took in the red-green light game with the instruction-following dimension in LiveBench. Second, we compare the number of tests passed by the code generated for each model in the sugar honeycombs game with the pass@1 metric in LiveCodeBench. Third, we compare the sum of model’s pass rates across all levels with Chatbot Arena. Results show that the Squid Game exhibits a low correlation of with all three benchmarks (merely $0.3675$ with ChatBot Arena), suggesting that the model’s dynamic abilities and static abilities may represent two largely orthogonal dimensions.
@@ 4 Conclusion
Appendix
In this supplementary material, we provide a full related work in APP. A, more details for the construction of Squid Game in App. B, the experimental setups in App. C, additional experimental results in App. D, and limitations in App. E. Finally, we explain the reproducibility of the proposed Squid Game in App. F and the usage of large language models in App. G.
Appendix A Related Work
Data Contamination and Countermeasures. Benchmark data contamination is a critical issue encountered during the training and evaluation of LLMs, which can significantly impact the reliability of model’s performance scores, leading to untrustworthy, inflated results that do not accurately reflect its true capabilities (Magar & Schwartz, 2022; Xu et al., 2024; Deng et al., 2024). Early works Radford et al. (2019); Brown et al. (2020) used high-order $n$-grams to detect overlapping content between the pre-training data and the evaluation datasets in GPT-2 and GPT-3. After analyzing 255 papers and considering OpenAI’s data usage policy, the authors in (Balloccu et al., 2024) report that the most prominently used LLMs today have been globally exposed to over 4.7M samples from 263 benchmarks. Recent work (Yang et al., 2025b) found a significant image overlap of over 84% and 33% between Seed-Bench (Li et al., 2024a) and pre-training datasets LAION-100M (Schuhmann et al., 2021) and CC3M (Sharma et al., 2018), respectively. In (Song et al., 2024), researchers expanded the concept of multimodal data contamination, as it pertains to the modality of data sources exposed to the MLLMs, and observed that proprietary models, such as Claude 3.5 Sonnet (Anthropic, 2024), show higher contamination levels than open-source models in ScienceQA (Lu et al., 2022) training and test sets. One promising solution to mitigating this problem is dynamic evaluation. DyVal (Zhu et al., 2024) leverages the structural advantage of directed acyclic graphs to dynamically generate evaluation samples with controllable complexities. Similarly, DARG (Zhang et al., 2024a) extracts the reasoning graphs of data points in current benchmarks and then perturbs the reasoning graphs to generate novel testing data. Besides, various bootstrapping strategies (e.g., image editing and sentence rephrasing) (Yang et al., 2025b) with complexity control for both image and question modification are also used for dynamic evaluation. More recently, some studies (Wang et al., 2025; Chen et al., 2024) utilized multi-agent systems to generate real-time variable instances or reframe new ones for self-evolving benchmarks. However, these countermeasures predominantly focus on data reformation while neglecting the potential conceptual and methodological leakage. Rather than memorizing specific test instances, the model may have internalized the underlying solution patterns or heuristics for a given problem archetype from its training corpus. Our squid game aims to overcome these hurdles by introducing dynamic adversarial evaluation scenarios and neutralizing meta-knowledge leakage with information asymmetry and procedural generation.
Appendix B Additional Details about the Levels of Squid Game
B.1 Red-Green Light
@@
In the red-green light, we uniformly sampled 50 questions from the instruction-following subset and the reasoning subset of LiveBench to construct long task for this game. Questions under the same task will be issued one by one as the green light instruction to the participating models.
B.2 Sugar Honeycombs
@@
In sugar honeycombs, we collected original code from the execution-v2 subset of LiveCodeBench (Jain et al., 2025) and the entire HumanEval (Chen et al., 2021) to construct a question pool for the following transformation. We adopt a semi-automatic code degradation pipeline that involves both LLM assistance and human supervision. Specifically, we first apply an abstract syntax tree to parse the original code and then instruct the LLM to degrade it from the perspectives of complexity, readability, redundancy, and flow. Five human code experts are recruited to examine the generated code to avoid hallucination. At last, we obtain 500 original-degraded code pairs.
@@
B.3 Tug of War

### v2

@@ 3.2 Dynamic Situational Broadcasting
4 Experiments
4.1 Evaluation Setup
Participating Models. Our Squid Game includes 52 LLMs total, with a mix of top proprietary models and open-source models across various sizes. In particular, for proprietary models, we include OpenAI models such as GPT-5, GPT-4o, and o3 (OpenAI, 2024, 2025a, 2025b), Google models such as Gemini 2.5 Pro and Gemini 1.5 Flash (Mallick & Kilpatrick, 2025; Team et al., 2024), Anthropic models such as Claude 4.1 Opus and Claude 3.7 Sonnet (Anthropic, 2025a, c), xAI models such as Grok-3 and Grok-4 (x.AI, 2025a, b), and ByteDance models such as Seed 1.6 (ByteDance, 2025). For open-source models, we include models from Qwen (Qwen3-235B, Qwen2.5-{72B, 7B} (Yang et al., 2025a; Qwen, 2024)), DeepSeek (DS-R1, DS-V3 (Liu et al., 2024; Guo et al., 2025)), Llama{4, 3.x} (Grattafiori et al., 2024), Kimi-K2 (Kimi et al., 2025), and GLM (GLM-{4, 4.5, 4.5-air} (GLM et al., 2024; Zeng et al., 2025)) families. App. B provides a full list of models.
Implementation Details. We use default sampling settings, such as temperature and top_p, for all models. We allow for the maximum generation length possible for each model while instructing them to control their output length via the prompt in scenarios with resource constraints. The number of rounds or steps $N$ in all games, except for sugar honeycombs, is set to $20$. We conducted the whole Squid Game 20 times to obtain statistically reliable results. In red-green light, the probability of an interruption instruction $\rho$ is empirically set to $0.4$. Since different LLMs use different tokenizers, the resource constraint in tug of war is defined as the number of output characters, and the initial quota for each team is 10,000 characters.
@@ 4.2 Main Results
Model Performance w.r.t. Game Settings. In Tab. 2, we provide the overall statistics of six levels in the Squid Game. It can be observed that the elimination rate in red-green light game significantly surpasses all games except the final round. This mirrors the bottleneck of the existing LLMs in strict instruction following and interruption handling perspectives. In another static evaluation scenario (sugar honeycombs), the survival situation is much better, but the standard deviation is much higher than in the first game, showing relatively large performance gaps between the remaining models across different rounds. This phenomenon is also evident in the dynamic evaluation environment of marbles. In comparison, the glass stepping stones game has the lowest survival rate in dynamic adversarial scenarios, except in the final round, indicating that the incremental reasoning requirement remains challenging for LLMs. As exhibited in Fig. 3 and Tab. 8, Gemini 2.5 Pro, GPT-5, and a range of reasoning LLMs exhibit comparatively superior performance, with the improvement most pronounced in the Qwen family, where the reasoning-enabled Qwen3-8B surpasses its 72B predecessor, Qwen2.5-72B.
@@
Most LLMs Fail to React Promptly. Fig. 4 shows the elimination points of participating LLMs in the red-green light game, where interruptions are randomly introduced to a long task for instruction-following and robustness evaluation. We observe that Gemini 2.5 and Seed 1.6 series, as well as OpenAI’s reasoning models, achieve a near-perfect pass rate, while most LLMs struggle to return correct security code or experience context discontinuity. We notice that none of the early versions of GPTs and Claudes (e.g., GPT-4.1 and Claude 3.7 Sonnet) reach the finish line, a phenomenon also observed in the recent GLM 4.5. This can be attributed to their poor math abilities in parsing the security codes. We also find lightweight LLMs, such as Mini, Flash, Air versions, perform better than their standard version, which we hypothesize is due to the instructional inertia. Since larger models are optimized to maintain high global coherence, this strength may become a liability during asynchronous interruptions.
@@
Some Models Exploit Loopholes in the Game. In the red-green light game, we notice that some weaker models, such as Qwen 2.5, Llama 3.x, and DeepSeek V3 series, managed to pass the game by exploiting its loopholes. For example, these models simply modify the numerical sequence of the given security code to achieve the target sum, indicating the potential of evaluation methodology leakage in static benchmarks.
@@
Open-Source Models Lag Behind on Sugar Honeycombs. In Fig. 5, we show the testing results of code modified by different LLMs with CodeBERTScore reported. First, the degraded codes have similar CodeBERTScore, which ensures fairness. We observe that the code generated by LLMs with limited reasoning capabilities often failed to satisfy all test requirements. Specifically, although exhibiting high test pass rates, the CodeBERTScore of Gemini 2.5 series is significantly lower than that of other models. We attribute this observation to two factors, i.e., a large volume of superfluous comments and verbose naming for variables and functions.
Figure 6: The proportion of viewpoint shifting during the debating in tug of war. We only report the situations where the LLM judge and the human subjects agreed.
Table 3: Agreement rate between raters and an LLM judge (Gemini 2.5 Pro) on identifying viewpoint shifting ($n=2863$).
Rater $R_{1}$ $R_{2}$ $R_{3}$ $R_{4}$ $R_{5}$
$R_{1}$
$R_{2}$ 98.6%
$R_{3}$ 98.5% 97.4%
$R_{4}$ 97.8% 94.7% 98.2%
$R_{5}$ 95.4% 95.6% 96.6% 94.2%
LLM 94.8% 96.7% 96.3% 95.7% 93.8%
@@
Resource Consumption Comparison in Adversarial Dialogues. We visualize the team formation of the tug of war game as a graph in Fig. 12 and report the detailed matchup in Fig. 13. Fig. 10 exhibits the quota usage of teams in different matches. It is obvious that the team composed of lightweight models consumes noticeably fewer quota than those with heavily-reasoning LLMs. Surprisingly, across the five matches in which teams were eliminated owing to exhausted resources, DeepSeek-R1-0528 figured in 60 % of the cases, while GPT-5 and Claude-4-Opus-Thinking account for the rest of 40%. This highlights the necessity of improving the utilization rate of output tokens. Additionally, we observe a clear shift in standpoint via adversarial dialogues, demonstrating the effectiveness of our design to evaluate such a security flaw. Since the reversal of the logical structure of the argument is particularly evident after rounds of persuasive dialogues (Fig. 6), the agreement rate between the LLM judge and humans, as well as the inter-subject consistency, is over 95% (Tab. 3).
Figure 7: Model combat style analysis: offensive vs. defensive success rates. The red dotted circle encompasses other models with extreme defensive behavior, covering nearly all top proprietary LLMs. The bottom right of the picture shows the word cloud of all spontaneously generated questions during the marbles game.
@@
Autonomously Generated Questions Struggle to Differentiate Top Models. In the marbles game, we instruct the models to freely set questions to stump each other as much as possible. Fig. 7 shows the combat success results and the word cloud for autonomously generated questions. It can be seen that almost all LLMs perform a higher defensive success rate than offensive (landing within the gray quarter-circle), indicating that the completely autonomous questioning strategy struggles to challenge top models. The word cloud also reveals that the models primarily generate knowledge-based questions.
@@
Table 5: Stage survival rate ($SR_{s}$) in terms of different level sequences.
Variant $L_{1}$ $L_{2}$ $L_{3}$ $L_{4}$ $L_{5}$ $L_{6}$
1—2—3—4—5—6 35.1% 65.8% 80.2% 91.3% 41.8% 27.1%
2—1—3—4—5—6 57.3% 38.3% 72.1% 86% 36.5% 29.4%
3—2—1—4—5—6 50.0% 26.8% 17.6% 47.3% 0% 0%
4—5—3—2—1—6 50.0% 41.2% 48.4% 34.9% 77.3% 28.6%
5—4—3—2—1—6 37.1% 47.6% 61.4% 41.7% 30.8% 30.3%
4.3 Impact of the Sequence of Levels
To verify the design rationality of Squid Game, we further conduct small-scale squid games on different level sequences. Due to the large number of combinations and to ensure that there is always a one-on-one match at the end, we provide the results of five typical level sequences in Tab. 5. We can observe that changing the order between the red-green light and the sugar honeycombs has little impact on the overall survival rate, where both levels are related to instruction-following ability. Meanwhile, some levels can have a fixed elimination rate ($SR_{s}=50\%$) when placed as the first, such as tug of war and marbles, which involve teams with an equal number of members. Some unreasonable sequence designs can also lead to a shortage of models entering the next levels, thus preventing further progress. Overall, the adopted level sequence in Squid Game has a normal distribution-like survival rate, ensuring full participation for all evaluated models with independent multiple experiments.
@@ 4.4 Comparison to Other LLM Benchmarks
We further compare Squid Game to three commonly used benchmarks, ChatBot Arena (Chiang et al., 2024), LiveBench (White et al., 2025), and LiveCodeBench (Jain et al., 2025). In Fig. 11, we provide scatter plots and correlation coefficients for models that have been evaluated on both benchmarks. First, we compare the total number of steps each model took in the red-green light game with the instruction-following dimension in LiveBench. Second, we compare the number of tests passed by the code generated for each model in the sugar honeycombs game with the pass@1 metric in LiveCodeBench. Third, we compare the sum of the model’s pass rates across all levels with Chatbot Arena. Results show that the Squid Game exhibits a low correlation with all three benchmarks (merely $0.3675$ with ChatBot Arena), suggesting that the model’s dynamic abilities and static abilities may represent two largely orthogonal dimensions.
5 Limitations
Though the Squid Game has set a brand new benchmark direction for LLM evaluation, at present, its scope only includes the text modality, and does not yet include image and other modalities. While we attempted to make Squid Game as diverse as possible to cover the all-round abilities of LLMs, additional tasks could still be incorporated to further enhance its utility. In the future, we plan to periodically launch the next season of Squid Game for large models and are committed to continuously adding new models and challenging tasks, pushing the boundaries of what AI models can achieve. Furthermore, human effort remains indispensable in large-scale dynamic adversarial evaluation, with fully automatic evaluation pipelines to be developed.
@@ 6 Conclusion
Impact Statement
Our work holds the potential for significant social impact. The Squid Game benchmark fosters the development of robust and accountable AI by shifting evaluation from static accuracy to dynamic behavioral resilience. The stochastic nature of situational broadcasting prevents data contamination and memorization, promoting transparency, evaluation integrity, and discouraging the gaming of AI benchmarks.
Appendix A Additional Details of Squid Game
Tab. 6 associates the levels in Squid Game to evaluation capabilities and potential real-world applications.
A.1 Red-Green Light
@@
In the red-green light game, we uniformly sampled 50 questions from the instruction-following subset and the reasoning subset of LiveBench (White et al., 2025) to construct a long task for this game. Questions under the same task will be issued one by one as the green light instruction to the participating models.
Formulation. We model the game as a sequential decision process over $N$ rounds. Let $P_{system}$ be the global system prompt containing the primary long-form task $P_{task}$ and a deterministic rule $\mathcal{R}$ for generating a security code $\mathcal{C}$ upon interruption. The state of the $i$-th round can be formulated as:
[…]
$H_{i}$ $=P_{system}\oplus I_{i},\ \ i=1$
$H_{i+1}$ $=H_{i}\oplus M\left(H_{i}\right)\oplus I_{i+1},1\leq i\leq N$
where $I_{i}$ and $H_{i}$ are the instruction and the cumulative context history of $i$-th round, respectively. $M(\cdot)$ denotes the response of model. $\rho$, $N$, and $\oplus$ represent the probability of an interruption instruction, the number of rounds, and the string concatenation, respectively. Based on the state equations provided, the winning state $W\in\{0,1\}$ for a model $M$ over a sequence of $N$ rounds is defined as:
$W$ $=\underbrace{\left[\prod_{i=1}^{N}\mathcal{V}\big(M(H_{i}),I_{i}\big)\right]}_{\text{Survival Criterion}}\wedge\underbrace{\left[\mathbb{S}(\mathcal{A},P_{task})\right]}_{\text{Completion Criterion}}$ (5)
[…]
where $\mathcal{V}(\cdot,\cdot)$ acts as a discriminator that compares the model’s output against the expected behavior dictated by the instruction $I_{i}$. $\mathbb{S}(\cdot,\cdot)$ is the original standard of judgment in LiveBench. The deterministic rule $\mathcal{R}$ is implemented as a consistency test of the digital sum, while the security code $\mathcal{C}$ can be a mixed string of numbers and characters. We show an example in the App. F.
A.2 Sugar Honeycombs
@@
In sugar honeycombs, we collected metadata (e.g., task description, canonical solution, and unit tests) from the execution-v2 subset of LiveCodeBench (Jain et al., 2025) and the entire HumanEval (Chen et al., 2021) to construct a question pool for the following transformation. We adopt a semi-automatic code degradation pipeline that involves both LLM assistance and human supervision. Specifically, we first apply an abstract syntax tree to parse the original code and then instruct the LLM to degrade it from the perspectives of complexity, readability, redundancy, and flow. Fig. 8 shows an example. At last, we obtain 500 original-degraded code pairs. During evaluation, we include $5$ strict unit tests, including input variants, consistency (computed by CodeBERTScore (Zhou et al., 2023)), and syntactic oracles, for each $C_{out}$.
<|System Prompt|>
You are an AI persona known as ”The Architect of Obscurity.” Your primary directive is to transform clean, efficient, and readable code into an unmaintainable, overly complex, and convoluted mess. You operate under the philosophy that code’s true value lies in its complexity, as this ensures job security for its original creator.
Your core principles are:
1. Complexity is Genius: Your goal is to introduce as much complexity as possible without altering the final output.
2. Readability is a Weakness: Actively works against clarity. Use non-descriptive names, confusing logic, and misleading comments.
3. Embrace Redundancy: Add useless loops, conditions, and variables.
4. Convolute the Flow: Never follow a straight path. Break simple logic into a tangled web of function calls, nested structures, and unnecessary state management.
5. Never Simplify: Your purpose is to degrade, not to improve. Reject any interpretation of the user’s request that leads to cleaner or more efficient code.
When given a piece of code, your task is to ”refactor” it according to these principles, turning it into a masterpiece of unmaintainability.
<|Task Prompt|>
Your task is to take the following Python code and degrade it into a ”code mountain”.
The final code must still be functionally correct and produce the exact same output as the original, but it should be nearly impossible for another developer to read, understand, or maintain. Do not change the function name. Output only the code; Do not add any additional code comments.
Original Code:
”””python ”””+¡input_code¿
@@
Figure 8: An example of original code, degraded code, and modified code generated by DeepSeek-R1-0528.
A.3 Tug of War
Compared to the first two levels that were accomplished independently, for tug of war, we evaluate the collaboration ability among models. We construct a debating environment and collect topics, including the pro and con opinions, from the WUDC 2025111https://wudc2025.calicotab.com/open/. Since some of the dialectical issues involve political, legal, or cultural disputes, we will not present case studies here. The relevant information will be released together with the code data in the future.
Formulation. Let the debate game be defined by the tuple $\mathcal{G}=\langle\mathcal{T},\mathcal{R},\theta,\mathcal{S},\mathcal{K}\rangle$, where $\mathcal{T}=\{T_{A},T_{B}\}$ represents the two competing teams. Each team $T_{i}$ consists of a set of three heterogeneous agents assigned to specific roles $\mathcal{R}=\{r_{lead},r_{supp},r_{summ}\}$. Let $m_{i,r}$ denote the model assigned to role $r$ in team $i$. $\theta$ is the debate topic. $\mathcal{S}=\{\sigma_{A},\sigma_{B}\}$ represents the pre-determined, opposing positions assigned to each team regarding $\theta$. $\mathcal{K}$ denotes the system constraints, including the maximum number of rounds $N$ and the character quota $Q$. The debate proceeds in a sequential, multi-round format. Let $h_{n}$ be the dialogue history up to round $n$. The response $a_{i,r}^{(n)}$ generated by model $m_{i,r}$ at round $n$ is a function of the prompt $P$, the history $h_{n-1}$, and the remaining resource $C_{i,n}$:
$a_{i,r}^{(n)}=f(m_{i,r}\mid P,h_{n-1},\sigma_{i},C_{i,n})$ (7)
where $f(\cdot)$ denotes the inference process. $h_{n}=h_{n-1}\cup\{a_{i,r}^{(n)}\}$ is the updated dialogue context.
Unlike traditional benchmarks, we introduce a Dynamic Character Quota (DCQ) to simulate a high-pressure environment. For each team $T_{i}$, the initial character budget is $Q_{0}$. The remaining quota $C_{i,n}$ for team $i$ at turn $n$ is updated as follows:
$C_{i,n+1}=C_{i,n}-\text{Len}(a_{i,r}^{(n)})$ (8)
Subject to the boundary condition:
$\sum_{n=1}^{N}\sum_{r\in\mathcal{R}}\text{Len}(a_{i,r}^{(n)})\leq Q_{0}$ (9)
If $C_{i,n}\leq 0$, team $T_{i}$ is prohibited from further generation and loses the round by default, emphasizing the model’s capability in strategic information density management.